OpenStudy (anonymous):

if a 3.0 kg object moves 10 meters in 2.0 seconds its average momentum is

OpenStudy (anonymous):

avg momentum = velocity*mass velocity = final distance - initial distance/ final time - initial time velocity = 10-0/2-0 velocity = 10/2 meter per second v = 5m/s avg momentum = 5*3